4(x+5)=3(3x-2)-2(x+2)
Found 2 solutions by Timnewman, ikleyn:
Answer by Timnewman(323)

Hi dear,
If you have,
4(x+5)=3(3x-2)-2(x+2)
and wish to find x,
clear the bracket as thus;
4x+20=9x-6-2x-4
4x+20=7x-10
collect like terms
7x-4x=20+10
3x=30
x=10

Answer by ikleyn(52778)

.
4(x+5) = 3(3x-2) - 2(x+2)

4x + 20 = 9x - 6 - 2x - 4

4x + 20 = 7x - 10

7x - 4x = 20 + 10

3x = 30

x = 10.

Check it yourself by substituting  x=10  into the original equations.